OpenStudy (anonymous):

what is an equation of the line, in point-slope form, that passes through the given point and had the given slope point: (8,-8); slope:3

OpenStudy (yttrium):

Just be reminded of this. Point-slope form is: \[y-y _{1} = m (x-x _{1})\] where m = slope x1 and y1 are the coordinates of the points on the line.

OpenStudy (anonymous):